OpenStudy (anonymous):

Which of the following are equal to the expression below? 2·√49 A) 98 B) 14 C) 196 D) √98 E) 49 F) √196

OpenStudy (anonymous):

b, f

OpenStudy (anonymous):

Lets first start by solving \[\sqrt{49}\] Do you know how to do that?

OpenStudy (anonymous):

Or maybe a better question is.... Do you know a number (hint one number between 1 - 9) that if you were to square it IE:\[5^2 = 25, 6^2=36... etc\] that you would get 49 ?

OpenStudy (anonymous):

Oh for √49 it is 7..

OpenStudy (anonymous):

Yes.......

OpenStudy (anonymous):

Correct! Another way of looking about how to solve square roots is you can rewrite them as following: \[\sqrt{49}=\sqrt{7*7}\] Because we are dealing with a square root (think of trying to group multiplies in groups of two) once we have this group of two, (so long as we're dealing only with multiplication) we can "pull" that number out. In this case, since we have two multiples of 7 under the square root, we can "pull" out one making it equal to 7. Now you can just do the rest of the math. 2 * 7 (i think... it kind of looks like 2-7 as well)